# Advancing TTP Analysis: Harnessing the Power of Encoder-Only and Decoder-Only Language Models with Retrieval Augmented Generation

### v1

@@ Abstract
Tactics, Techniques, and Procedures (TTPs) outline the methods attackers use to exploit vulnerabilities. The interpretation of TTPs in the MITRE ATT&CK framework can be challenging for cybersecurity practitioners due to presumed expertise, complex dependencies, and inherent ambiguity. Meanwhile, advancements with Large Language Models (LLMs) have led to recent surge in studies exploring its uses in cybersecurity operations. This leads us to question how well encoder-only (e.g., RoBERTa) and decoder-only (e.g., GPT-3.5) LLMs can comprehend and summarize TTPs to inform analysts of the intended purposes (i.e., tactics) of a cyberattack procedure. The state-of-the-art LLMs have shown to be prone to hallucination by providing inaccurate information, which is problematic in critical domains like cybersecurity. Therefore, we propose the use of Retrieval Augmented Generation (RAG) techniques to extract relevant contexts for each cyberattack procedure for decoder-only LLMs (without fine-tuning). We further contrast such approach against supervised fine-tuning (SFT) of encoder-only LLMs. Our results reveal that both the direct-use of decoder-only LLMs (i.e., its pre-trained knowledge) and the SFT of encoder-only LLMs offer inaccurate interpretation of cyberattack procedures. Significant improvements are shown when RAG is used for decoder-only LLMs, particularly when directly relevant context is found. This study further sheds insights on the limitations and capabilities of using RAG for LLMs in interpreting TTPs.
Introduction
The evolving changes of cybersecurity vulnerabilities presents a significant challenge for incident and threat analysis. Cyber Threat Intelligence (CTI) reports focus on gathering, analyzing, and interpreting data about potential cyber attacks. CTI aims to understand and anticipate the Tactics, Techniques, and Procedures (TTPs) of cyber adversaries. The process of deciphering and understanding the various cyberattack procedures (which is constantly updated in reference models like the MITRE ATT&CK [^1] https://attack.mitre.org/), requires extensive expertise and effort. The ATT&CK framework serves as a crucial tool in this regard, offering in-depth insights into adversary behaviors and techniques in targeting systems. The ATT&CK framework utilizes a 3/4-tier Tactic-Technique/Sub-technique-Procedure hierarchy to explain the how and why behind attackers’ exploitation of vulnerabilities. However, the complexity and potential ambiguity of these descriptions can make them challenging to interpret, leading to differing conclusions among security analysts from the same description. For example, consider the following attack procedure description:
@@ Introduction
The ATT&CK framework associates three tactics with this procedure, namely, ‘Privilege-Escalation’, ‘Defense-Evasion’, and ‘Persistence’. It is not obvious that the DLL Search Order Hijacking technique will lead to any of the three, not to mention all three adversarial intents. In fact, some might argue that this description is more aligned with the ‘Execution’ tactic, as the adversary tries to execute its malicious payload. Therefore, we ask: Can LLMs interpret and map procedure descriptions to specific ATT&CK tactics? Note that we recognize the inherent ambiguity of ATT&CK TTP mapping. This in fact motivated us to study the capabilities and limitations of LLMs under such real-world imperfections.
@@
The evolution of Large Language Models (LLMs), such as the OpenAI’s GPT-3.5 model [^2] https://platform.openai.com/docs/models, has led to substantial progress in NLP tasks (Min et al. 2023; Zhao et al. 2023). This progress is mainly due to the LLM’s enhanced semantic understanding and scalability. There have also been works that process TTPs by leveraging LLMs (Sauerwein and Pfohl 2022; Husari et al. 2017; You et al. 2022; Rani et al. 2023; Alves, Geraldo Filho, and Gonçalves 2022; Fayyazi and Yang 2023). However, none of these works compared the use of encoder-only LLMs (e.g., RoBERTa) and decoder-only LLMs (e.g., GPT-3.5) while leveraging Retrieval Augmented Generation (RAG) techniques to analyze the descriptions of attack procedures.
@@
The encoder-only LLMs excel in tasks like sequence classification and entity recognition by processing input data to generate contextual representations, while decoder-only models adept at generating coherent, contextually relevant text, making them ideal for language generation tasks. One major problem of LLMs is the tendency to hallucinate. This issue involves the model’s ability to produce responses that are misleading or completely fabricated, yet presented confidently. To address this issue, RAG techniques are proposed to extract the most relevant documents and feed those as additional information into the input of LLMs. It is used to shift the model’s attention to relevant information by first retrieving documents that are likely to contain the answer and then using this information to generate a response. This method is beneficial as it is not only limited to model’s temporal knowledge but also plays a crucial role in reducing hallucinations by grounding responses from selected resources.
@@
Recognizing the potentially different uses and benefits for encoder-only and decoder-only LLMs for cybersecurity operations, this study sets out to compare the two in their effectiveness to interpret cyberattack procedures and map them to ATT&CK tactics. Specifically, we examine three sets of approaches: 1) direct use of the decoder-only LLMs (GPT-3.5-Turbo), 2) supervised fine-tuning (SFT) of encoder-only LLMs (RoBERTa (Liu et al. 2019) and SecureBERT (Aghaei et al. 2022)) with ATT&CK descriptions, and 3) the use of RAG on decoder-only LLMs. Note that in all three approaches, we avoid explicitly training of LLMs to map the procedure descriptions to MITRE ATT&CK tactics. This is meant to test LLM’s interpretation capability instead of building a traditional classifier model. In a way, we hope to mimic, to the extent possible, how LLMs perform if the procedure description does not exist on the Internet.
@@
To the best of our knowledge, this is the first formal study to compare these approaches to interpret cyberattack procedure descriptions. By experimenting using the latest TTP descriptions extracted from MITRE ATT&CK (release v14.1 as of Oct. 31, 2023), we present the benefits and limitations of the two LLM classes and suggest future research directions. Our main contributions are as follows:
@@
  • We analyze and compare the performance in SFT of smaller encoder-only LLMs vs. RAG-enhanced of larger decoder-only LLMs for interpreting TTPs.
  • We demonstrate a significant improvement in interpreting TTPs for decoder-only LLMs when the right pieces of information (via RAG) is provided.
  • We study the different retrieval methodologies to provide relevant context of procedure descriptions and show the limitations of RAG techniques when they fail to retrieve relevant information.
@@
  • We demonstrate that while decoder-only LLMs maintain strong ’recall’, they lack ’precision’, and we elaborate the need to new advances in increasing the ‘precision’ in LLMs without compromising ‘recall’.
  • We demonstrate the current capabilities of using RAG with decoder-only LLMs by providing specific examples, and offer insights into their lack of ‘interpretability’.
Related Works
Large Language Models
@@
The advancements in Large Language Models (LLMs) are largely attributed to the adoption of the Transformer architecture (Vaswani et al. 2017). Central to this architecture are the encoder and decoder components, which have been utilized distinctively in various LLMs.
@@
Encoder-Only: One of the pioneering introductions is BERT (Devlin et al. 2019), a pre-trained encoder-only LLM based on transformers. It leverages the Transformer encoder to achieve a deep understanding of language context and nuances. BERT’s significance lies in its bidirectional training, which enables a comprehensive understanding of language structure. Moreover, RoBERTa (Liu et al. 2019), extended BERT’s capabilities. It optimized hyperparameters, eliminated the Next Sentence Prediction task, and enhanced training with larger mini-batches and longer sequences. Another notable encoder-only model is SecureBERT (Aghaei et al. 2022), which is based on RoBERTa but specifically fine-tuned with cybersecurity data.
@@
Decoder-Only Models: The landscape of LLMs has been further transformed by decoder-only models, characterized by their extensive training datasets and significant computational resources. OpenAI’s GPT-3.5 and Meta’s LLAMA-2333https://ai.meta.com/llama are leading examples in this category, both encompassing billions of parameters. These models have not only shown the ability to scale but also their proficiency in generating human-like text and recognizing intricate patterns across various fields. The effectiveness and versatility of these models in handling diverse and complex tasks have been extensively documented in the literature, as described in (Min et al. 2023; Zhao et al. 2023). However, there remains a question on how capable these models are in interpreting TTPs.
LLM for TTP interpretations
In the context of processing TTP descriptions, several works have been proposed by using NLP techniques. (Sauerwein and Pfohl 2022; Husari et al. 2017; Kim, Kim et al. 2022).
@@
More recently, the following works have utilized encoder-only LLMs (e.g., BERT) for classifying TTPs. You et al. (2022) introduced the TIM framework, using continuous description sets and SentenceBERT embeddings, along with TTP-specific features like IPs and URLs, to categorize TTPs based on five techniques and one tactic of the ATT&CK framework. Alves, Filho, and Gonçalves (2022) tested various BERT model variants with different hyperparameters to find the the best model for TTP classification. Rani et al. (2023) developed TTPHunter, a tool that uses BERT and RoBERTa models with a linear classifier for extracting TTPs from unstructured text. Orbinato et al. (2022) conducted an experimental study to compare traditional ML and DL models (e.g., SVM vs. SecureBERT) for mapping TTPs from Cyber Threat Intelligence texts. The authors found SecureBERT to be the most effective, however, they did not provide a comparison between SecureBERT and its foundational RoBERTa model.
Furthermore, our previous work (Fayyazi and Yang 2023) compared the encoder-only vs the direct use of decoder-only LLMs to process and classify higher-level tactic and technique descriptions. The encoder-only experiments focused on traditional classifier training where training and testing sets are drawn from the same set of data. The decoder-only approach did not integrate with Retrieval Augmented Generation. The current work significantly extend our previous findings to treat harder-to-interpret attack procedure descriptions and with more advanced LLM approaches without explicitly mapping the testing data to tactic labels.
Retrieval Augmented Generation (RAG)
@@
Despite the strong capabilities of state-of-the-art LLMs in instruction-following tasks (e.g., question-answering), these models have shown to be prone to hallucination in their generated responses. Their reliance on pre-trained knowledge (i.e., constrained by the training data’s temporal scope) poses limitations, particularly in domains where accuracy is paramount like cybersecurity. Fine-tuning with fresh data can be expensive and still prone to hallucination. Therefore, to adapt to new information and to provide factual knowledge, RAG was introduced (Borgeaud et al. 2022).
@@
In RAG, relevant pieces of information is stored in a vector database. When a query is presented, RAG selects the most relevant data from this database, which is then incorporated into the prompt of LLMs before they generate a response. This approach offers two key advantages: first, it allows the LLM to incorporate current information without needing constant fine-tuning; second, it enhances the clarity and relevance of the responses by providing contextually grounded information. Essentially, RAG equips LLMs with the ability to gain a focused understanding of the specific context of a query. This approach improves the clarity and precision of the responses by directing the model’s attention towards the most relevant context retrieved. There have been works that used RAG in NLP tasks (Borgeaud et al. 2022; Mallen et al. 2023; Ram et al. 2023; Al Ghadban et al. 2023). In this work, we propose the use of RAG to provide more information about an attack procedure to the LLM and focus the LLM’s attention towards the text retrieved through RAG.
Methodology & Experimental Design
@@
We design this study to compare: 1) supervised fine-tuning (SFT) of encoder-only LLMs with labeled ATT&CK technique/sub-technique descriptions, 2) direct use of decoder-only LLMs (i.e., pre-trained knowledge), and 3) use of RAG for decoder-only LLMs with relevant URLs retrieved by finding the most similar attack procedures. Furthermore, we consider an ideal (yet unrealistic) case where the exact URL of each procedure is used for RAG to give the upper bound performance. This experimental design enables us to assess how well LLMs interpret cyberattack procedure descriptions and map them to the corresponding tactic(s). This is a challenging process due to multiple reasons: 1) there are no publicly available datasets that map procedure descriptions into ATT&CK tactic(s), 2) each procedure can map to multiple ATT&CK tactics that LLMs need to consider, 3) a well-structured prompt is needed for LLMs to generate desired outputs, and 4) RAG needs to retrieve the most relevant information. We discuss next our methodology and design of experiments to tackle these challenges.
Figure 1: The process of supervised multi-label fine-tuning of encoder-only LLMs with high-level ATT&CK descriptions (excluding ‘procedures’) – adopted from (Fayyazi and Yang 2023).
Datasets
@@
To conduct our experiments, we gathered data from the MITRE ATT&CK framework. We chose ATT&CK due to its broad adoption in industry SIEM tools and associated detailed cybersecurity descriptions. We curated the descriptions of enterprise tactics, techniques, and sub-techniques along with the mapping to their corresponding tactic(s) for fine-tuning the encoder-only models. In total, we obtained 639 descriptions from the MITRE ATT&CK framework. Some descriptions have 2,3, or even 4 different tactics. Note that the use of 639 labeled ATT&CK descriptions for supervised fine-tuning of encoder-only LLMs reflects the real-world scenario where only limited labeled data would exist. Moreover, to test the performance of both encoder-only and decoder-only LLMs, we crawled all the enterprise procedure examples that describe a attack methodology from the ATT&CK framework. We remove the procedure descriptions that contain any of the 14 MITRE ATT&CK tactics within the descriptions to prevent potential biases or shortcuts of the LLMs. In total, we gathered 9,532 procedure descriptions along with their corresponding URLs (for RAG). Note that the procedure descriptions are not used for SFT of encoder-only LLMs. Figure 2 shows the pair-wise overlap across tactics for attack procedure descriptions. As can be seen, ‘Persistence’, ‘Privilege-Escalation’ and ‘Defense-Evasion’ tactics have higher overlaps with the others. We will share the curated data along with our source code as artifacts for the research community.
Figure 2: The pair-wise overlap of tactics on entire ATT&CK procedure descriptions.
Supervised Fine-Tuning of Encoder-Only LLMs

### v3

# Advancing TTP Analysis: Harnessing the Power of Large Language Models with Retrieval Augmented Generation

@@ Abstract
Tactics, Techniques, and Procedures (TTPs) outline the methods attackers use to exploit vulnerabilities. The interpretation of TTPs in the MITRE ATT&CK framework can be challenging for cybersecurity practitioners due to presumed expertise and complex dependencies. Meanwhile, advancements with Large Language Models (LLMs) have led to recent surge in studies exploring its uses in cybersecurity operations. It is, however, unclear how LLMs can be used in an efficient and proper way to provide accurate responses for critical domains such as cybersecurity. This leads us to investigate how to better use two types of LLMs: small-scale encoder-only (e.g., RoBERTa) and larger decoder-only (e.g., GPT-3.5) LLMs to comprehend and summarize TTPs with the intended purposes (i.e., tactics) of a cyberattack procedure. This work studies and compares the uses of supervised fine-tuning (SFT) of encoder-only LLMs vs. Retrieval Augmented Generation (RAG) for decoder-only LLMs (without fine-tuning). Both SFT and RAG techniques presumably enhance the LLMs with relevant contexts for each cyberattack procedure. Our studies show decoder-only LLMs with RAG achieves better performance than encoder-only models with SFT, particularly when directly relevant context is extracted by RAG. The decoder-only results could suffer low ‘Precision’ while achieving high ‘Recall’. Our findings further highlight a counter-intuitive observation that more generic prompts tend to yield better predictions of cyberattack tactics than those that are more specifically tailored.  [^1] Code available at: https://github.com/RezzFayyazi/TTP-LLM
1 Introduction
The evolving changes of cybersecurity vulnerabilities presents a significant challenge for incident and threat analysis. Security analysts utilize Cyber Threat Intelligence (CTI) to gather, analyze, and interpret the Tactics, Techniques, and Procedures (TTPs) of adversaries. The process of understanding the various cyberattack procedures (which is constantly updated in reference models like MITRE ATT&CK MITRE ), requires extensive expertise and effort. The ATT&CK framework utilizes a 3/4-tier Tactic-Technique/Sub-technique-Procedure hierarchy to explain the how and why behind attackers’ exploitation of vulnerabilities. However, the complexity and potential ambiguity of these descriptions can make them challenging to interpret, leading to differing conclusions among analysts from the same description. For example, consider the following attack procedure description:
@@ 1 Introduction
The ATT&CK framework associates three tactics with this procedure, namely, ‘Privilege Escalation’, ‘Defense Evasion’, and ‘Persistence’. It is not obvious that the DLL Search Order Hijacking technique will lead to any of the three, not to mention all three adversarial intents. In fact, some might argue that this description is more aligned with the ‘Execution’ tactic, as the adversary tries to execute its malicious payload.
@@
Therefore, in this work, we delve into the potential of Large Language Models (LLMs) to interpret and map procedure descriptions into specific ATT&CK tactics. Note that we recognize the inherent imperfection of ATT&CK TTP mapping, and this, in fact, motivated us to study the capabilities and limitations of LLMs under such real-world constraints. The evolution of LLMs, such as the OpenAI’s GPT-3.5 model OpenAI , has led to substantial progress in NLP tasks Min et al. (2023); Zhao et al. (2023). This progress is mainly due to the LLM’s enhanced semantic understanding and scalability. There have also been works that process TTPs by leveraging LLMs Rani et al. (2023); Alves et al. (2022); Orbinato et al. (2022). However, none of these works compared the use of encoder-only LLMs (e.g., RoBERTa) and decoder-only LLMs (e.g., GPT-3.5) while leveraging Retrieval Augmented Generation (RAG) techniques to analyze the descriptions of attack procedures.
@@
The encoder-only LLMs (e-LLM) excel in tasks like sequence classification by processing input data to generate contextual representations, while decoder-only models (d-LLM) are adept at generating coherent, contextually relevant text, making them ideal for language generation tasks. One major problem of d-LLMs is the tendency to hallucinate Huang et al. (2023); Rawte et al. (2023). This issue involves the model’s ability to produce responses that are misleading or completely fabricated, yet presented confidently. To address this issue, RAG techniques are proposed to extract the most relevant documents and feed those as additional information into part of the LLM prompts. It is used to shift the model’s attention to relevant information by first retrieving texts that are likely to contain relevant information to generate a more validated response. Note that we are not fine-tuning the d-LLMs, as such a process requires significant computational resources. Instead, we consider RAG to be a potentially efficient technique to provide d-LLMs with cybersecurity context.
@@
Recognizing the potentially different uses and benefits for e-LLMs and d-LLMs for cybersecurity operations, this study sets out to compare the supervised fine-tuning (SFT) of e-LLMs and RAG with d-LLMs to interpret TTPs in their effectiveness to interpret cyberattack procedures (assessed by predicting ATT&CK tactics). Specifically, we examine four sets of approaches: 1) direct use of the d-LLMs (GPT-3.5), 2) SFT of e-LLMs (RoBERTa Liu et al. (2019) and SecureBERT Aghaei et al. (2022)) with ATT&CK descriptions, and 3) the use of RAG on d-LLMs and 4) analysis of d-LLM capabilities with more specific vs. more generic prompts to predict cyberattack tactics.
@@
To the best of our knowledge, this is the first formal study to compare these approaches to interpret cyberattack procedure descriptions. We aim to derive ‘practical’ approaches that are reasonable for cybersecurity practitioners to adopt. By experimenting using the TTP descriptions extracted from MITRE ATT&CK (release v14.1 as of Oct. 31, 2023), we present the benefits and limitations of the two LLM architectures and suggest future research directions to enhance d-LLM’s use in cyber-ops. Our contributions includes:
@@
  • We analyze and compare the performance in SFT of smaller e-LLMs vs. RAG-enhanced of larger d-LLMs for interpreting TTPs.
  • We demonstrate d-LLMs achieve a superior performance in interpreting TTPs when directly relevant information is extracted via RAG.
@@
  • We show that while d-LLMs maintain strong ’Recall’, they lack ’Precision’, and we elaborate the need for new advances in increasing the ‘precision’ in d-LLMs without compromising ‘recall’.
  • We present preliminary results showing that the use of generic prompts seemingly provide more informative and accurate responses than the use of specific prompts.
2 Related Works
2.1 Large Language Models
@@
The advancements in Large Language Models are primarily due to the adoption of the Transformer architecture Vaswani et al. (2017). Central to this architecture are the encoder and decoder components, which is utilized distinctively in various LLMs.
@@
Encoder-Only: One of the pioneering introductions is BERT Devlin et al. (2019), a pre-trained bidirectional e-LLM based on Transformers. It leverages the Transformer encoder to achieve a deep understanding of language context and nuances. Moreover, RoBERTa Liu et al. (2019), extended BERT’s capabilities with optimized hyperparameters and enhanced training with larger mini-batches and longer sequences. Another notable e-LLM is SecureBERT Aghaei et al. (2022), which is based on RoBERTa but specifically fine-tuned with cybersecurity data.
@@
Decoder-Only Models: The landscape of LLMs has been further transformed by d-LLMs, characterized by their extensive training datasets and significant computational resources. OpenAI’s GPT-3.5 OpenAI and Meta’s LLAMA-2 Touvron et al. (2023) are leading examples in this category, both with proficiency in generating human-like text and recognizing intricate patterns across various fields. The effectiveness and versatility of these models in handling diverse and complex tasks have been extensively documented in the literature, as described in Min et al. (2023); Zhao et al. (2023). However, there remains a question on how capable these models are in interpreting TTPs.
2.2 LLM for TTP interpretations
@@
In the context of processing TTP descriptions, several works have been proposed to use NLP techniques. Sauerwein & Pfohl (2022); Husari et al. (2017); Kim et al. (2022). More recently, the following works have utilized e-LLMs (e.g., BERT) for classifying TTPs. Alves et al. Alves et al. (2022) tested various BERT model variants with different hyperparameters to find the the best model for TTP classification. Rani et al. Rani et al. (2023) developed TTPHunter, a tool that uses BERT and RoBERTa models with a linear classifier for extracting TTPs from unstructured text. Orbinato et al. Orbinato et al. (2022) conducted an experimental study to compare traditional ML and DL models (e.g., SVM vs. SecureBERT) for mapping TTPs from CTI texts. The authors found SecureBERT to be the most effective, however, they did not provide a comparison between SecureBERT and its foundational RoBERTa model.
Our preliminary study Fayyazi & Yang (2023) compared e-LLMs vs. the direct use of d-LLMs (i.e., prompt engineering) to process and classify higher-level tactic and technique descriptions (but not procedures). In that study, we fine-tuned the e-LLM classifier by drawing from the pool of tactic and technique descriptions, and the d-LLM approach did not include with RAG. This paper significantly expands our prior work’s findings to treat harder-to-interpret attack procedure descriptions. It adopts a more realistic setting where procedure descriptions are not used for fine-tuning and incorporates RAG techniques.
2.3 Retrieval Augmented Generation (RAG)
@@
Despite the proficiency of d-LLMs in instruction-following tasks (e.g., question-answering), these models have shown to be prone to hallucination in their generated responses Huang et al. (2023); Rawte et al. (2023). Their reliance on pre-trained knowledge (i.e., constrained by the training data’s temporal scope) poses limitations. This limitation is critical in fields requiring high precision, such as cybersecurity. Therefore, to adapt to new information and to provide factual knowledge, RAG was introduced Borgeaud et al. (2022).
@@
RAG typically utilizes a vector database to store up-to-date information. Upon receiving a query, RAG selects relevant data from this database to enrich the LLM’s prompt, thereby ensuring responses are both current and contextually grounded without constant fine-tuning. There have been works that used RAG in NLP tasks Borgeaud et al. (2022); Mallen et al. (2023); Ram et al. (2023); Al Ghadban et al. (2023). In this work, we propose the use of RAG to provide more information about an attack procedure to the LLM and shift the LLM’s attention towards the text retrieved through RAG.
3 Methodology & Experimental Design
@@
We design this study to compare: 1) supervised fine-tuning (SFT) of e-LLMs with labeled ATT&CK technique/sub-technique descriptions, 2) direct use of d-LLMs (i.e., pre-trained knowledge), 3) use of RAG for d-LLMs with relevant URLs retrieved by finding the most similar attack procedures, 4) analysis of d-LLM capabilities in interpreting attack procedures using a specific vs. generic prompt. This experimental design enables us to assess how well LLMs interpret cyberattack procedure descriptions and map them to the corresponding tactic(s). This is a challenging process due to multiple reasons: 1) there are no publicly available datasets that map procedure descriptions into ATT&CK tactic(s), 2) each procedure can map to multiple ATT&CK tactics that LLMs need to consider, 3) a well-structured prompt is needed for LLMs to generate desired outputs, and 4) RAG needs to retrieve the most relevant information. We outline our methodology and experimental design to tackle these challenges.
3.1 Datasets
@@
To conduct our experiments, we gathered data from the MITRE ATT&CK framework. We chose ATT&CK due to its broad adoption in industry SIEM tools and associated detailed cybersecurity descriptions. We curated the descriptions of enterprise tactics, techniques, and sub-techniques along with the mapping to their corresponding tactic(s) for fine-tuning the e-LLMs. We do not fine-tune the targeted d-LLMs (GPT-3.5); instead, we use RAG for d-LLMs to compared to the fine-tuned e-LLM. In total, we obtained 639 descriptions from the ATT&CK framework for fine-tuning e-LLMs. Some descriptions have 2,3, or even 4 different tactics. Note that the small volume of 639 labeled ATT&CK descriptions reflects the real-world scenario where only limited labeled data would exist for targeted fine-tuning. To test the performance of both e-LLMs and d-LLMs, we crawled all the enterprise procedure examples that describe an attack methodology from the ATT&CK framework. We removed the procedure descriptions that contain any of the 14 MITRE ATT&CK tactics within the descriptions to prevent potential biases or shortcuts for the LLMs. In total, we gathered 9,532 procedure descriptions along with their corresponding URLs (for RAG). Note that the procedure descriptions are not used for SFT of e-LLMs. Figure 1 shows the pair-wise overlap across tactics for attack procedure descriptions. As can be seen, ‘Persistence’, ‘Privilege-Escalation’ and ‘Defense-Evasion’ tactics have higher overlaps with the others. We will share the curated data along with our source code as artifacts for the research community.
Figure 1: The pair-wise overlap of tactics on ATT&CK procedure descriptions.
3.2 Supervised Fine-Tuning of Encoder-Only LLMs
Figure 2: The process of supervised multi-label fine-tuning of e-LLMs with high-level ATT&CK descriptions (excluding ‘procedures’) – adopted from Fayyazi & Yang (2023).